# The surprising efficiency of temporal difference learning for rare event prediction

### v2

@@ 3 Experiments
Effect of the initial distribution. So far, in our experiments, we have chosen $\mu$ to be the uniform distribution on $D$. A natural alternative strategy is to harvest many short trajectories from a single, much longer trajectory. Under ergodicity assumptions on $P$, these short trajectories are then approximately drawn from the invariant distribution of $P$. In Fig. 4, we present the relative asymptotic variance and the empirical estimates of the relative MSE of the mean first passage time and the committor, with $P$ being the same as in (8), and $\mu$ chosen to be the invariant distribution $p$ in (9) conditioned within $D$. For $\tau=1$ and $n=20,40,$ and $80$, remarkably, the maximum relative asymptotic variances of the LSTD estimators grow exponentially with $n$. With $M=10n^{3}$, the empirical relative MSE estimators with $30000$ independent copies of the LSTD estimators agree well with the relative asymptotic variance for $n=20$ and $40$. For $n=80$, the empirical linear system (3) fails to define the LSTD estimators with high probability, and the empirical error is undefined. These results should be contrasted with the much smaller errors shown in the middle panels of Figs. 1 and 2, corresponding to uniformly chosen initial conditions. The results indicate that the choice of $\mu$ can also have a significant impact on the performance of LSTD for prediction of rare events.
@@ 4 Rare event assumptions and upper bound
For some constant $c>0$, independent of $n$, the minorizing graph $G$ is connected.
When $P$ is irreducible, as assumed in Theorem 1, and $n$ is fixed, we can always find a $c>0$ so that $G$ is connected. Assumption 2 emphasizes the independence of $n$, asserting that the transition probabilities above the threshold $c$ can, on their own, preserve the connectivity of the graph while $n$ increases. Recall that the Markov chain studied in Section 3 satifies $P(i,i\pm 1)\geq 1/(2(1+e))$, implying that Assumption 2 is satisfied with the choice $c=1/(2(1+e))$.
@@ Appendix B Properties of the Markov chain studied in Section 3
We can find an explicity formula for $u$ by solving the linear system
@@
$u(k)$ $=2\left(\sum_{i=1}^{n-1}\frac{p(i)+p(i+1)}{p(i)p(i+1)}\right)^{-1}\left[\left(\sum_{i=1}^{k-1}\frac{p(i)+p(i+1)}{p(i)p(i+1)}\right)\left(\sum_{i=k}^{n-1}\frac{p(i)+p(i+1)}{p(i)p(i+1)}\left(\sum_{l=2}^{i}p(\ell)\right)\right)\right.$
$\left.-\left(\sum_{i=k}^{n-1}\frac{p(i)+p(i+1)}{p(i)p(i+1)}\right)\left(\sum_{i=2}^{k-1}\frac{p(i)+p(i+1)}{p(i)p(i+1)}\left(\sum_{l=2}^{i}p(\ell)\right)\right)\right].$ (15)
@@
$\left[\left(\sum_{i=1}^{\frac{n+1}{2}-1}\frac{p(i)+p(i+1)}{p(i)p(i+1)}\right)\left(\sum_{i=1}^{\frac{n-1}{2}}\frac{p(i)+p(i+1)}{p(i)p(i+1)}\left(\sum_{\ell=2}^{\frac{n+1}{2}}p(\ell)+\sum_{l=2}^{i}p(\ell)\right)\right)\right.$
@@
$=2\left(\sum_{i=1}^{n-1}\frac{p(i)+p(i+1)}{p(i)p(i+1)}\right)^{-1}\left[\left(\sum_{i=1}^{\frac{n-1}{2}}\frac{p(i)+p(i+1)}{p(i)p(i+1)}\right)^{2}\left(\sum_{l=2}^{\frac{n+1}{2}}p(\ell)\right)\right]$
@@ Appendix C Proofs of main results
Let $\tilde{M}^{t}$ denote a matrix containing only the (normalized) numerators of $\tilde{S}^{t}-S^{t}$, that is, $\tilde{M}^{t}(k,\ell)=\frac{1}{M}\sum_{j=1}^{M}\mathbf{1}_{\{k\}}(X_{0}^{j})\left(\mathbf{1}_{\{\ell\}}(X_{t\wedge T^{j}}^{j})-S^{t}(k,\ell)\right)$. Let $\tilde{M}^{1:\tau}$ represent a long vector formed by flattening out each matrix $M^{t}$ in row major order and gluing up in time order. Then, since trajectories are independent, $\tilde{M}^{1:\tau}$ satisfies a central limit theorem as
@@
Before explicitly calculating this covariance, we deal with the denominators first. Let $\tilde{\mu}$ be the vector containing all the (normalized) denominators, that is, $\tilde{\mu}(k)=\frac{1}{M}M_{k}$.
Then, again because of independence of trajectories, we have a weak law of large numbers as
@@
As for $\Sigma_{\tilde{M}}^{1:\tau}$, the key fact is that $\mathbf{1}_{\{(k,\ell)\}}$ for different $k$ are independent. Therefore,
Combining (23), (24), (26) and (27), by Slutsky’s theorem, we obtain a central limit theorem for $\tilde{S}$ as
@@
We define hitting times counting from $t=0$. That is, $\hat{T}_{k}=\min\{t\geq 0:X_{t}=k\}$. This differs from $T_{k}$ because $\hat{T}_{k}$ includes $t=0$. Similarly, we define $\hat{T}^{\tau}_{k}$ and more generally $\hat{T}_{A}$ and $\hat{T}_{A}^{\tau}$ for any subset $A\subseteq[n]$ as the counterparts of $T_{k}^{\tau}$, $T_{A}$, and $T_{A}^{\tau}$ but counted from $t=0$. But notice that we have defined $T$ and $T^{\tau}$ to be counted from $t=0$. In all hitting time definitions, when the set over which the minimum is taken is empty, the hitting time is defined to +$\infty$+ ∞.
@@
$S_{j}(k,\ell)=\begin{cases}S(k,\ell)&k\neq j,k,\ell\in[n]\\ \delta_{k\ell}&k=j,l\in[n].\end{cases}$ (42)
@@
$\begin{split}\mathbf{P}_{\ell}\left[T_{j}^{\tau,j}<T_{\Delta}^{{\tau,j}}\right]&=Q(\ell,\ell)\mathbf{P}_{\ell}\left[T_{j}^{\tau,j}<T_{\Delta}^{\tau,j}\right]+Q(\ell,k)\mathbf{P}_{k}\left[T_{j}^{\tau,j}<T_{\Delta}^{\tau,j}\right]+Q(\ell,j),\\ \mathbf{P}_{k}\left[T_{j}^{\tau,j}<T_{\Delta}^{\tau,j}\right]&=Q(k,k)\mathbf{P}_{k}\left[T_{j}^{\tau,j}<T_{\Delta}^{\tau,j}\right]+Q(k,\ell)\mathbf{P}_{l}\left[T_{j}^{\tau,j}<T_{\Delta}^{\tau,j}\right]+Q(k,j).\end{split}$ (45)
@@
$\frac{{Q}(k,\ell)\overline{Q}(\ell,j)+{Q}(k,j)}{\mathbf{P}_{k}\left[T_{\ell}^{\tau}<T_{k}^{\tau}\wedge T_{\Delta}^{{\tau}}\right]}\geq\frac{\overline{Q}(k,\ell)\overline{Q}(\ell,j)+\overline{Q}(k,j)}{\overline{Q}(k,\ell)+\overline{Q}(k,j)}\geq\overline{Q}(\ell,j).$
@@
Now we deal with the exceptional cases. When $\ell\in D$ but $j=k$ or $j=\ell$, the same conclusion holds with minor modifications of the intermediate steps. Specifically, take $j=\ell$ (so that $j\in D$ and $\Delta=D^{\textrm{c}}$) as an example. Then $Z_{t}\in\{\ell,k,\Delta\}$ records transitions of $Y_{t}^{\tau,j}$ among these three states, with $Q$ denote its transition probabilities.
$\mathbf{P}_{\ell}\left[\hat{T}_{j}^{\tau,j}\leq T^{\tau,j}\right]-\mathbf{P}_{k}\left[\hat{T}_{j}^{\tau,j}\leq T^{\tau,j}\right]=1-\overline{Q}(k,j)=\overline{Q}(k,\Delta),$
and $\mathbf{P}_{k}\left[T_{\ell}^{\tau}<T_{k}^{\tau}\wedge T^{\tau}\right]=Q(k,j)$, $\mathbf{P}_{k}\left[T_{j}<T\right]=\overline{Q}(k,j)$. Therefore,
@@
Then the following steps are the same as (48) in the general case.
@@
As for the expectation following $e_{i}^{\mathsf{T}}(I-S_{D}^{\tau})^{-1}e_{k}$ in (5), we note from the fourth equality in (37) that
@@
By the Bellman equation (2), $\mathbf{E}\left(\sum_{t=1}^{\tau-1}R_{D}(X_{t\wedge T})+u(X_{\tau\wedge T})\,\Big{|}\,X_{0}=k,X_{1\wedge T}\right)=u(X_{1\wedge T})$. And the first term can be simplified as
$\begin{split}&\mathbf{E}_{k}\left[\mathbf{Var}\left(\sum_{t=1}^{\tau-1}R_{D}(X_{t\wedge T})+u(X_{\tau\wedge T})\,\Big{|}\,X_{0}=k,X_{1\wedge T}\right)\right]\\ =&\mathbf{E}_{k}\left[\mathbf{Var}\left(\sum_{t=2}^{\tau-1}R_{D}(X_{t\wedge T})+u(X_{\tau\wedge T})\,\Big{|}\,X_{0}=k,X_{1\wedge T}\right)\right],\end{split}$
and then we can further condition it on $X_{2\wedge T}$ and perform a similar decomposition as in (50). This conditioning can be done inductively for $X_{t\wedge T}$ up to $t=\tau$. As a result of this inductive conditioning, we obtain

### v3

@@ 3 Experiments
Effect of the initial distribution. So far, in our experiments, we have chosen $\mu$ to be the uniform distribution on $D$. A natural alternative strategy is to harvest many short trajectories from a single, much longer trajectory. Under ergodicity assumptions on $P$, these short trajectories are then approximately drawn from the invariant distribution of $P$. In Fig. 4, we present the relative asymptotic variance and the empirical estimates of the relative MSE of the mean first passage time and the committor, with $P$ being the same as in (8), and $\mu$ chosen to be the invariant distribution $p$ in (9) conditioned within $D$. For $\tau=1$ and $n=20,40,$ and $80$, the maximum relative asymptotic variances of the LSTD estimators grow exponentially with $n$. With $M=10n^{3}$, the empirical relative MSE estimators with $30000$ independent copies of the LSTD estimators agree well with the relative asymptotic variance for $n=20$ and $40$. For $n=80$, the empirical linear system (3) fails to define the LSTD estimators with high probability, and the empirical error is undefined. These results should be contrasted with the much smaller errors shown in the middle panels of Figs. 1 and 2, corresponding to uniformly chosen initial conditions. We expect that the results in Fig. 4 would be somewhat different if, instead of initializing trajectories independently from the conditional invariant distribution, we had harvested correlated initial conditions from a single ergodic trajectory. Nonetheless, the results indicate that the choice of $\mu$ can have a significant impact on the performance of LSTD for prediction of rare events.
@@ 4 Rare event assumptions and upper bound
For some constant $c>0$ independent of $n$, any two states $i\in D$ and $j\in[n]$ are connected by a path in $G$.
When $P$ is irreducible, as assumed in Theorem 1, and $n$ is fixed, we can always find a $c>0$ so that any $i\in D$ and $j\in[n]$ are connected by a path in $G$ for $\tau=1$. Assumption 2 emphasizes the independence of $n$, asserting that the transition probabilities above the threshold $c$ can, on their own, preserve the connectivity of the graph while $n$ increases. Recall that the Markov chain studied in Section 3 satifies $P(i,i\pm 1)\geq 1/(2(1+e))$, implying that Assumption 2 is satisfied with the choice $c=1/(2(1+e))$.
@@ Appendix B Properties of the Markov chain studied in Section 3
We can find an explicit formula for $u$ by solving the linear system
@@
$u(k)$ $=2\left(\sum_{i=1}^{n-1}\frac{p(i)+p(i+1)}{p(i)p(i+1)}\right)^{-1}\left[\left(\sum_{i=1}^{k-1}\frac{p(i)+p(i+1)}{p(i)p(i+1)}\right)\left(\sum_{i=k}^{n-1}\frac{p(i)+p(i+1)}{p(i)p(i+1)}\left(\sum_{\ell=2}^{i}p(\ell)\right)\right)\right.$
$\left.-\left(\sum_{i=k}^{n-1}\frac{p(i)+p(i+1)}{p(i)p(i+1)}\right)\left(\sum_{i=2}^{k-1}\frac{p(i)+p(i+1)}{p(i)p(i+1)}\left(\sum_{\ell=2}^{i}p(\ell)\right)\right)\right].$ (15)
@@
$\left[\left(\sum_{i=1}^{\frac{n+1}{2}-1}\frac{p(i)+p(i+1)}{p(i)p(i+1)}\right)\left(\sum_{i=1}^{\frac{n-1}{2}}\frac{p(i)+p(i+1)}{p(i)p(i+1)}\left(\sum_{\ell=2}^{\frac{n+1}{2}}p(\ell)+\sum_{\ell=2}^{i}p(\ell)\right)\right)\right.$
@@
$=2\left(\sum_{i=1}^{n-1}\frac{p(i)+p(i+1)}{p(i)p(i+1)}\right)^{-1}\left[\left(\sum_{i=1}^{\frac{n-1}{2}}\frac{p(i)+p(i+1)}{p(i)p(i+1)}\right)^{2}\left(\sum_{\ell=2}^{\frac{n+1}{2}}p(\ell)\right)\right]$
@@ Appendix C Proofs of main results
Let $\tilde{M}^{t}$ denote a matrix containing only the (normalized) numerators of $\tilde{S}^{t}-S^{t}$, that is, $\tilde{M}^{t}(k,\ell)=\frac{1}{M}\sum_{j=1}^{M}\mathbf{1}_{\{k\}}(X_{0}^{j})\left(\mathbf{1}_{\{\ell\}}(X_{t\wedge T^{j}}^{j})-S^{t}(k,\ell)\right)$. Let $\tilde{M}^{1:\tau}$ represent a long vector formed by flattening out each matrix $\tilde{M}^{t}$ with $1\leq t\leq\tau$ in row major order and gluing up in time order. Then, since trajectories are independent, $\tilde{M}^{1:\tau}$ satisfies a central limit theorem as
@@
Since $\mathbf{1}_{\{k\}}(X_{0})$ for different $k$ can not be nonzero at the same time, we immediately obtain
Now we deal with the denominators. Let $\tilde{\mu}$ be the vector containing all the (normalized) denominators, that is, $\tilde{\mu}(k)=\frac{1}{M}M_{k}$. Then, again because of independence of trajectories, we have a weak law of large numbers as
@@
Combining (23), (24), (26), and (27) , by Slutsky’s theorem, we obtain a central limit theorem for $\tilde{S}$ as
@@
We define additional hitting times counting from $t=0$. That is, $\hat{T}_{k}=\min\{t\geq 0:X_{t}=k\}$. This differs from $T_{k}$ because $\hat{T}_{k}$ includes $t=0$. Similarly, we define $\hat{T}^{\tau}_{k}$ and more generally $\hat{T}_{A}$ and $\hat{T}_{A}^{\tau}$ for any subset $A\subseteq[n]$ as the counterparts of $T_{k}^{\tau}$, $T_{A}$, and $T_{A}^{\tau}$ but counted from $t=0$. But notice that we have defined $T$ and $T^{\tau}$ to be counted from $t=0$. In all hitting time definitions, when the set over which the minimum is taken is empty, the hitting time is defined to be +$\infty$+ ∞.
@@
$S_{j}(k,\ell)=\begin{cases}S(k,\ell)&k\neq j,k,\ell\in[n]\\ \delta_{k\ell}&k=j,\ell\in[n].\end{cases}$ (42)
@@
$\begin{split}\mathbf{P}_{\ell}\left[T_{j}^{\tau,j}<T_{\Delta}^{{\tau,j}}\right]&=Q(\ell,\ell)\mathbf{P}_{\ell}\left[T_{j}^{\tau,j}<T_{\Delta}^{\tau,j}\right]+Q(\ell,k)\mathbf{P}_{k}\left[T_{j}^{\tau,j}<T_{\Delta}^{\tau,j}\right]+Q(\ell,j),\\ \mathbf{P}_{k}\left[T_{j}^{\tau,j}<T_{\Delta}^{\tau,j}\right]&=Q(k,k)\mathbf{P}_{k}\left[T_{j}^{\tau,j}<T_{\Delta}^{\tau,j}\right]+Q(k,\ell)\mathbf{P}_{\ell}\left[T_{j}^{\tau,j}<T_{\Delta}^{\tau,j}\right]+Q(k,j).\end{split}$ (45)
@@
$\frac{{Q}(k,\ell)\overline{Q}(\ell,j)+{Q}(k,j)}{\mathbf{P}_{k}\left[T_{\ell}^{\tau}<T_{k}^{\tau}\wedge T^{{\tau}}\right]}\geq\frac{\overline{Q}(k,\ell)\overline{Q}(\ell,j)+\overline{Q}(k,j)}{\overline{Q}(k,\ell)+\overline{Q}(k,j)}\geq\overline{Q}(\ell,j).$
@@
Now we deal with the exceptional cases. When $\ell\in D$ but $j=k$ or $j=\ell$, the same conclusion holds with minor modifications of the intermediate steps. Specifically, take $j=\ell$ (so that $j\in D$ and $\Delta=D^{\textrm{c}}$) as an example. Then $Z_{t}\in\{\ell,k,\Delta\}$ records transitions of $Y_{t}^{\tau,j}$ among these three states. With $Q$ denoting its transition probabilities.
$\mathbf{P}_{\ell}\left[\hat{T}_{j}^{\tau,j}\leq T^{\tau,j}\right]-\mathbf{P}_{k}\left[\hat{T}_{j}^{\tau,j}\leq T^{\tau,j}\right]=1-\overline{Q}(k,j)=1-\overline{Q}(k,\ell)=\overline{Q}(k,\Delta),$
and $\mathbf{P}_{k}\left[T_{\ell}^{\tau}<T_{k}^{\tau}\wedge T^{\tau}\right]=Q(k,\ell)$, $\mathbf{P}_{k}\left[T_{j}<T\right]=\overline{Q}(k,\ell)$. Therefore,
@@
We also note that $\mathbf{P}_{k}\left[T^{\tau}<T_{k}^{\tau}\right]\geq Q(k,\Delta)$. Then the following steps are the same as (47) and (48) in the general case.
@@
As for the expectation following $e_{i}^{\mathsf{T}}(I-S_{D}^{\tau})^{-1}e_{k}$ in (5), we note from the Bellman equation (2) that
@@
By the definition of $u$, $\mathbf{E}\left(\sum_{t=1}^{\tau-1}R_{D}(X_{t\wedge T})+u(X_{\tau\wedge T})\,\Big{|}\,X_{0}=k,X_{1\wedge T}\right)=u(X_{1\wedge T})$, which helps to simplify the second term. Meanwhile, the first term can be simplified as
$\begin{split}&\mathbf{E}_{k}\left[\mathbf{Var}\left(\sum_{t=1}^{\tau-1}R_{D}(X_{t\wedge T})+u(X_{\tau\wedge T})\,\Big{|}\,X_{0}=k,X_{1\wedge T}\right)\right]\\ =&\mathbf{E}_{k}\left[\mathbf{Var}\left(\sum_{t=2}^{\tau-1}R_{D}(X_{t\wedge T})+u(X_{\tau\wedge T})\,\Big{|}\,X_{0}=k,X_{1\wedge T}\right)\right].\end{split}$
We can further condition it on $X_{2\wedge T}$ and perform a similar decomposition as in (50), which can be done inductively for $X_{t\wedge T}$ up to $t=\tau$. As a result of this inductive conditioning, we obtain